OpenStudy (anonymous):

1) F(x)=4^xlog(small=>)7(x)] 2) f(x)=[Inx]^4 3)f(x)=2x^(4x) 4)f(x)=In(sqrt(6x+7)(7x-3))

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Okay, what's difficult?

OpenStudy (anonymous):

Logarithmic Derivatives!

OpenStudy (anonymous):

\[ (\ln(F(x)))' = \frac{F'(x)}{F(x)} \]

OpenStudy (anonymous):

You want to take the logarithm of both sides, simplify things, and then implicitly differentiate.

OpenStudy (anonymous):

how do i do that? :O

OpenStudy (anonymous):

Is this correct?\[ F(x)=4^x\log_7(x) \]

OpenStudy (anonymous):

yes, thars right!

OpenStudy (anonymous):

There is no need to do this one with logarithmic differentiation, it'd be easier to do product rule.

OpenStudy (anonymous):

yea i was told by someone that but my teacher want me to do it by using Logarithmic Derivatives rules

OpenStudy (anonymous):

could you teach me please?

OpenStudy (anonymous):

\[ \ln[F(x)]=\ln[4^x\log_7(x)] = x\ln(4)+\ln[\log_7(x)] \]Differentiate both sides: \[ \frac{F'(x)}{F(x)} = \ln(4) + \frac{\frac{1}{x\ln(7)}}{\log_7(x)} \]

OpenStudy (anonymous):

Put back in the value for \(F(x)\)\[ \frac{F'(x)}{4^x\log_7(x)} = \ln(4) + \frac{\frac{1}{x\ln(7)}}{\log_7(x)} \]Simplify a bit: \[ F'(x) = 4^x\log_7(x)\ln(4) + \frac{4^x}{x\ln(7)} \]

OpenStudy (anonymous):

thats really cool how you solve the problem easily. the last sentence is the answer ?

OpenStudy (anonymous):

I showed every step...

OpenStudy (anonymous):

thank you. !

OpenStudy (anonymous):

how do the next question?

OpenStudy (anonymous):

One question per help.

OpenStudy (anonymous):

you wont help me?